What is the electric force from the rod on the point charge Q? Suppose that Q-13.1 nC and L = 4.07 cm. Break the rod up into four equal pieces and use superposition to find the force 2 mC 4L
Answer #1

to the right end of the rodis, The distance from the charge di = 14 +(32)* = VIOL The distance from the charge dx = +(L)* = VFrom the figure, 2018 – 2008 - cose, - ਝ = cose, -ਤੇੜ The Inear charge density is, The electric field due to charged rod alon

The electric force along the x-axis is, F, = QE, = (13.1x10°)(-1.0618x10°)=-13.91 N The electric force along the y-axis is, F